OpenStudy (anonymous):

How do I find the ratio of the volume to the surface area for the rectangular prism: Total SA = 10x^2 + 40x, V= 4x^2+12x-16?

OpenStudy (anonymous):

The equation you gave of V has dimmension of m square??

OpenStudy (anonymous):

V is surface area of what??

OpenStudy (anonymous):

SA = 10x(x+4) V =4 (x+4)(x-1) v/sa = 4(x-1) / 10x = 2(x-1) / 5x

OpenStudy (anonymous):

at jimmyrep, V is not here volume you can see by dimmension

OpenStudy (anonymous):

you are right - i hadnt noticed that!